OpenStudy (anonymous):

6cos^2 theta + 5 cos theta -4 = 0

OpenStudy (anonymous):

u can factorise it \[(2\cos \theta -1)(3\cos \theta+4)=0\]

OpenStudy (anonymous):

to find the exact solution set?

OpenStudy (anonymous):

so u will get \[\theta= \cos^{-1} (1/2) \] and \[\theta = \cos^{-1} (-4/3)\]

OpenStudy (anonymous):

but how is this done?

OpenStudy (anonymous):

after factorised. note that 2(cos theta )-1=0 and 3(cos theta) +4=0 rearrange them and solve both equations to get theta.

OpenStudy (anonymous):

ahh, i get it. thanks(:

OpenStudy (anonymous):

no prob